# Learned representation-guided diffusion models for large-image generation

### v1

@@ Abstract
To synthesize high-fidelity samples, diffusion models typically require auxiliary data to guide the generation process. However, it is impractical to procure the painstaking patch-level annotation effort required in specialized domains like histopathology and satellite imagery; it is often performed by domain experts and involves hundreds of millions of patches. Modern-day self-supervised learning (SSL) representations encode rich semantic and visual information. In this paper, we posit that such representations are expressive enough to act as proxies to fine-grained human labels. We introduce a novel approach that trains diffusion models conditioned on embeddings from SSL. Our diffusion models successfully project these features back to high-quality histopathology and remote sensing images. In addition, we construct larger images by assembling spatially consistent patches inferred from SSL embeddings, preserving long-range dependencies. Augmenting real data by generating variations of real images improves downstream classifier accuracy for patch-level and larger, image-scale classification tasks. Our models are effective even on datasets not encountered during training, demonstrating their robustness and generalizability. Generating images from learned embeddings is agnostic to the source of the embeddings. The SSL embeddings used to generate a large image can either be extracted from a reference image, or sampled from an auxiliary model conditioned on any related modality (e.g. class labels, text, genomic data). As proof of concept, we introduce the text-to-large image synthesis paradigm where we successfully synthesize large pathology and satellite images out of text descriptions.
@@ 1 Introduction
Modern-day self-supervised learning (SSL) representations [5] encode rich semantic and visual information. Features from trained self-supervised models serve as compact image representations and are widely used to perform discriminative downstream tasks successfully [6, 46, 11]. We hypothesize that such SSL representations are already expressive enough to act as proxies to fine-grained human labels. If this is true, these representations should be able to condition the training of effective diffusion models in these domains. In this novel approach, we utilize self-supervised feature extractors as image annotators; these features provide necessary per-image conditioning signals at the highest resolutions, required for diffusion model training.
@@ 2 Related work
Self-Supervised Learning Self-supervised learning refers to both discriminative [13, 2] and distillation [5] approaches that aim to learn useful representations of the data without any supervision. Diffusion models have been shown to match the performance of SSL approaches [48]. In this work, we are mainly interested in self-supervised learning for histopathology. Notable recent developments include the Hierarchical Image Pyramid Transformer (HIPT) [6], which utilizes the inherent hierarchical structure of Whole Slide Images (WSIs), CTransPath [46], a hybrid CNN and multi-scale Swin Transformer model, and iBOT [11], a masked image modeling method. These models excel as patch-level feature extractors for WSIs by leveraging the unique structure of the large-image data.
@@ 3.2 Large image generation
$F_{i,j}:\mathcal{J}\rightarrow I,\quad\lambda_{i,j}:\mathcal{Z}\rightarrow\mathcal{Y}$ (4)
where, for the case of large-image generation, we set $F_{i,j}$ to be a crop of the large image, centered at $i,j$, and $\lambda_{i,j}$ the conditioning $y_{i,j}\in\mathbb{R}^{d}$ at $i,j$. Since we only have conditioning vectors at the centers of the patches, we can use spatial interpolation algorithms to implement $\lambda$. This assumes that the interpolant $y_{i,j}$ is a valid conditioning vector for the diffusion process, which we validate experimentally. Then, $\Psi$ is defined as
$\Psi(J_{t}\mid z)=\underset{J\in\mathcal{J}}{\arg\min}\sum_{i,j}\lVert F_{i,j}(J)-\Phi(F_{i,j}(J_{t})\mid\lambda_{i,j}(z))\rVert^{2}$ (5)
@@
$\Psi(J_{t}\mid z)=\sum_{i,j}\frac{F_{i,j}^{-1}(\mathbf{1})}{\sum\limits_{k,l}F_{k,l}^{-1}(\mathbf{1})}\otimes F_{i,j}^{-1}(\Phi(F_{i,j}(J_{t})\mid\lambda_{i,j}(z)))$ (6)
where $\mathbf{1}$ denotes a patch image where all values are set to 1.
We are able to generate images larger than the ones produced by the patch diffusion model. At the same time, we control what each patch looks like, which is crucial in maintaining the semantic integrity of the larger image. The self-supervised conditions can capture the variations between neighboring patches necessary for producing realistic results. Generating large images with coarser conditioning, such as global text prompts [50], would lead to uniform texture regions with visible discontinuities between them (see supplementary).
@@ 3.3 Controllable large-image synthesis
Since there is no available dataset of paired large images and captions, we resort to pre-trained multi-modal models, such as Quilt [18], CLIP [35] and BLIP [22], to provide the text conditioning. We construct training sets by i) extracting self-supervised embeddings from training-set large images and ii) pairing them with multi-modal image embeddings or machine-generated captions. During inference, we first sample SSL embeddings from the learned distribution $p(z\mid c)$, and then utilize our patch diffusion models to synthesize a large image.
@@ 4.1 Datasets
Figure 3: Synthetic patches (256 $\times$ 256) from our models, and the corresponding reference real images used to generate them. The SSL-guided LDM replicates the semantics of the reference patch.
Figure 4: Large images (1024 $\times$ 1024) from our models, and the corresponding reference real images used to generate them. We preserve the global arrangement of the semantics defined in the reference image.
@@ 4.2 Implementation details
We train our models on $6$ NVIDIA RTX 8000 GPUs, with a batch size of $100$ per GPU, utilizing code and pre-trained checkpoints from LDM [38]. We set the learning rate at $10^{-4}$ with a warmup of 10,000 steps. We apply DDIM with $50$ steps and a guidance scale of $1.75$ for both patch sampling and large-image generation. When generating large images we apply the patch diffusion model with a stride $>1$ depending on the desired target quality. More details are found in the supplementary.
@@ 4.3 Image quality results
In Fig. 3 and 4, we present synthetic patches and large images from our TCGA-BRCA and NAIP models, along with the corresponding references from which the self-supervised embeddings were extracted. We evaluate our method’s per-patch and large-image generation quality by computing FID scores [14] using the Clean-FID implementation [32]. We generate 10,000 patches ($256\times 256$) and 3,000 large images ($1024\times 1024$) from diffusion models trained on the TCGA-BRCA, TCGA-CRC, and NAIP datasets. Since our generative model requires a self-supervised conditioning vector $y$ (or multiple vectors $z$) for each synthetic image (or large image), we randomly sample embeddings from reference images in the training set to generate images for evaluation.
For patches, we measure FID against the real images (“Vanilla FID”). For large images, we follow the evaluation strategy of MultiDiffusion [3] and use FID to compare the distribution of 256 $\times$ 256 crops from synthesized large images to that of real image patches of the same size (“Crop FID”). We also measure FID directly between the large images and ground truth data using CLIP [35] (“CLIP FID”).
To evaluate the similarity between the reference and generated large images, we resize from $1024\times 1024$ to $256\times 256$ and use an SSL model to extract embeddings. We compute the cosine similarity between the paired embeddings (“Embedding Similarity”).
Patch-level quality: Our models achieve low patch FID scores across all datasets (Tab. 1). For TCGA-BRCA patches, our “Vanilla FID” of 6.98 is comparable to the current state-of-the-art [50] ($7.64$ at 10×10$\times$10 ×). We attain similar, low FID scores for the smaller CRC and NAIP datasets.
@@
Table 1: FID scores for our generated patch and large images. Our patch-level BRCA model is on par with SoTA [50] (7.64 at 10 $\times$). “CLIP FID” and “Embedding Similarity” demonstrate our large images’ realism and contextual accuracy.
Synthetic large image quality: We present the “Crop FID”, “CLIP FID” and “Embedding Similarity” results for large images in Table 1. The “Crop FID” of the large images is comparable to the “Vanilla FID” on the BRCA and CRC datasets, showing that patches from synthesized large images have similar semantic content to the ground truth patches. We attribute the worse “Crop FID” for the NAIP model to the limited number of samples available for both SSL and diffusion model training. The “Crop FID” is consistently higher than the patch-level FID, which is expected as the large-image generation framework only approximates the distribution of the large images and does not have access to the true conditioning at every location.
Our low “CLIP FID” scores indicate that the generated large images are similar to real images when resized to $224\times 224$ pixels. This indicates that our SSL-guided large-image generation successfully retains the larger-scale semantic arrangements found in real data. Unfortunately, for NAIP, our model is not as good at synthesizing high-frequency details, which explains the large discrepancy between “Crop FID” and “CLIP FID”. Additionally, when comparing the “CLIP FID” of large images synthesized by the 20×20$\times$20 × BRCA model and resized to 5×5$\times$5 × to images from a model trained directly on 5×5$\times$5 × data, we see that the difference is minimal (7.43 vs. 6.64).
We evaluate the contextual similarity between synthetic and reference large images for the BRCA and CRC data. We compute cosine similarity between large images using CTransPath embeddings. Our BRCA and CRC models demonstrate high “Embedding Similarity” scores of $0.924$ and $0.938$, respectively, reflecting our framework’s effectiveness in preserving the integrity of context and key features on the large-image scale.
@@ 5 Image Augmentation Experiments
As shown in Fig. 3 and 4, apart from visual fidelity, the synthetic images preserve the intricate characteristics of the reference images, both on the patch and large-image scale. These characteristics include the nuanced textural elements, histological staining, and cell structure. This correspondence between real and synthetic images demonstrates the detailed and varied information captured by the self-supervised embeddings used as conditioning. In conjunction with the high “Embedding Similarity”, it justifies using variations of images generated from SSL embeddings for patch and large-image level data augmentation.
@@
Table 2: The inclusion of synthetic data consistently enhances AUC across various MIL architectures and BRCA tasks. The dataset contains 1000 real images, so “10% + synthetic” indicates training with 100 real and 100 synthetic WSIs, with the remainder used for testing.
@@
The results, presented in Table 2(a), reveal a notable improvement in classifier performance, from $78\%$ to $83\%$. This improvement again confirms the quality of our synthetic data while also highlighting our approach’s versatility. Despite being trained on $256\times 256$ pixel patches from TCGA-BRCA, the model generalizes to produce high-quality large images on a completely different dataset. We attribute this generalization capability to the expressiveness of the SSL features and the potency of the diffusion model to accurately portray them in images. While our model does not reach the current state-of-the-art accuracy of $87\%$, achieved by an ensemble approach [7], the simplicity and ability to integrate with other models make for a noteworthy contribution.
@@
Table 3: Our data augmentations provide notable improvements for the BACH (a) and CRC-VAL-HE (b) datasets. Notably, in both cases, the diffusion training data does not overlap with the augmented dataset.
@@
We train an ImageNet pre-trained ResNet-50 [12] network on three splits: real images only, synthetic images only, and a combination of both. Evaluation on the CRC-VAL-HE-7K test set demonstrates a significant performance increase when synthetic data is introduced, with classifier accuracy rising from 93.8% to 96.27% as presented in Table 2(b). We match the current SoTA [20], which used an ensemble of deep models, with a model agnostic approach. As in the previous experiment, the diffusion model, now trained on the significantly smaller TCGA-CRC data, can also effectively synthesize images from a completely different dataset by only controlling the self-supervised conditioning (details in supplementary).
@@ 7 Combining self-supervised embeddings with diffusion
We further evaluate our patch-generating models by posing the question does the diffusion model learn more about the data than the self-supervised learning model? We hypothesize that by combining the pre-trained self-supervised embedding with the denoising task we can improve the learned representations of the data, leading to better performance in downstream tasks, which are performed with features from self-supervised learning. To validate this hypothesis, we utilize the trained diffusion model as a feature extractor and apply a Multiple-instance learning (MIL) approach for the slide-level classification task of subtyping Breast Cancer.
We validate the hypothesis that the SSL-conditioned diffusion model fuses the information from the two separate tasks leveraged during training: the self-supervised and denoising tasks. We use the diffusion model as a feature extractor and apply MIL algorithms for slide-level classification of breast cancer subtypes. For each patch in a whole-slide image (WSI) we first extract the self-supervised embeddings, which are then used as conditioning to obtain features from the denoiser’s U-Net bottleneck layer at a fixed timestep $t=50$.
@@ Acknowledgements
The reported research was partially supported by NCI awards 5U24CA215109, 1R21CA258493-01A1, UH3CA225021, NSF grants IIS-2123920, IIS-2212046 and Stony Brook Profund 2022 seed funding. We would also like to thank Rajarsi Gupta for his valuable feedback.
@@ 9 Annotation costs
We asked a pathologist to annotate patches from TCGA-BRCA to estimate the cost of detailed per-patch annotation for the entirety of the dataset. We presented the 20×20$\times$20 × magnification patches of Fig. 6 and requested them to “write a brief description for each of the following patches”. An expert pathologist required approximately 5-10 seconds to identify features and describe the patches. Therefore, for the entire 15M patches of TCGA-BRCA, it would take $\approx 40000$ hours to provide full per-patch annotations. This training dataset is small compared to the volume of data used in large studies, e.g. 10k whole slide images or approximately 10×10$\times$10 × the number of TCGA-BRCA data. Employing expert pathologists to annotate these vast amounts of data is prohibitively expensive and, therefore, practically infeasible at the scale at which we want to apply these models.

### v2

@@ Abstract
To synthesize high-fidelity samples, diffusion models typically require auxiliary data to guide the generation process. However, it is impractical to procure the painstaking patch-level annotation effort required in specialized domains like histopathology and satellite imagery; it is often performed by domain experts and involves hundreds of millions of patches. Modern-day self-supervised learning (SSL) representations encode rich semantic and visual information. In this paper, we posit that such representations are expressive enough to act as proxies to fine-grained human labels. We introduce a novel approach that trains diffusion models conditioned on embeddings from SSL. Our diffusion models successfully project these features back to high-quality histopathology and remote sensing images. In addition, we construct larger images by assembling spatially consistent patches inferred from SSL embeddings, preserving long-range dependencies. Augmenting real data by generating variations of real images improves downstream classifier accuracy for patch-level and larger, image-scale classification tasks. Our models are effective even on datasets not encountered during training, demonstrating their robustness and generalizability. Generating images from learned embeddings is agnostic to the source of the embeddings. The SSL embeddings used to generate a large image can either be extracted from a reference image, or sampled from an auxiliary model conditioned on any related modality (e.g. class labels, text, genomic data). As proof of concept, we introduce the text-to-large image synthesis paradigm where we successfully synthesize large pathology and satellite images out of text descriptions.  [^1] Code is available at this link
@@ 1 Introduction
Modern-day self-supervised learning (SSL) representations [5] encode rich semantic and visual information. Features from trained self-supervised models serve as compact image representations and are widely used to perform discriminative downstream tasks successfully [6, 46, 11], proving that these compressed representations indeed encode useful semantic information about the images. We hypothesize that such SSL representations are already expressive enough to act as proxies to fine-grained human labels. If this is true, these representations should be able to condition the training of effective diffusion models in these domains. In this novel approach, we utilize self-supervised feature extractors as image annotators; these features provide necessary per-image conditioning signals at the highest resolutions, required for diffusion model training.
@@ 2 Related work
Self-Supervised Learning Self-supervised learning (SSL) refers to both discriminative [13, 2] and distillation [5] approaches that aim to learn representations of the data without supervision. In this work, we are mainly interested in self-supervised learning for histopathology. Notable recent developments include the Hierarchical Image Pyramid Transformer (HIPT) [6], which utilizes the inherent hierarchical structure of Whole Slide Images (WSIs), CTransPath [46], a hybrid CNN and multi-scale Swin Transformer model, and iBOT [11], a masked image modeling method. These models excel as patch-level feature extractors for WSIs by leveraging the unique structure of the large-image data and capture the important semantic information that we require for conditioning a generative model.
@@ 3.2 Large image generation
$F^{\text{proj}}_{i,j}:\mathcal{J}\rightarrow I,\quad\lambda_{i,j}:\mathcal{Z}\rightarrow\mathcal{Y}$ (4)
where, for the case of large-image generation, we set $F^{\text{proj}}_{i,j}$ to be a crop (projection) of the large image, centered at $i,j$, and $\lambda_{i,j}$ the conditioning $y_{i,j}\in\mathbb{R}^{d}$ at $i,j$. Since we only have conditioning vectors at the centers of the patches, we can use spatial interpolation algorithms to implement $\lambda$. This assumes that the interpolant $y_{i,j}$ is a valid conditioning vector for the diffusion process, which we validate experimentally. Then, $\Psi$ is defined as
$\Psi(J_{t}\mid z)=\underset{J\in\mathcal{J}}{\arg\min}\sum_{i,j}\lVert F^{\text{proj}}_{i,j}(J)-\Phi(F^{\text{proj}}_{i,j}(J_{t})\mid\lambda_{i,j}(z))\rVert^{2}$ (5)
@@
$\Psi(J_{t}\mid z)=\sum_{i,j}\frac{F^{\text{unproj}}_{i,j}(\mathbf{1})}{\sum\limits_{k,l}F^{\text{unproj}}_{k,l}(\mathbf{1})}\otimes F^{\text{unproj}}_{i,j}(\Phi(F^{\text{proj}}_{i,j}(J_{t})\mid\lambda_{i,j}(z)))$ (6)
where $\mathbf{1}$ denotes a patch image where all values are set to 1 and $F^{\text{unproj}}_{i,j}$ is the inverse mapping of pixels from the crop centered at $i,j$ back to the large image.
We are able to generate images larger than the ones produced by the patch diffusion model. At the same time, we control what each patch looks like, which is crucial in maintaining the semantic integrity of the larger image. The self-supervised conditions can capture the variations between neighboring patches necessary for producing realistic results. Generating large images with coarser conditioning, such as global text prompts [49], would lead to uniform texture regions (see supplementary).
@@ 3.3 Controllable large-image synthesis
Since there is no available dataset of paired large images and captions, we resort to pre-trained multi-modal models, such as Quilt [18], CLIP [35] and BLIP [22], to provide the text conditioning. We construct training sets by extracting self-supervised embeddings from training-set large images and pairing them with multi-modal image embeddings or generated captions. During inference, we first sample SSL embeddings from the learned distribution $p(z\mid c)$, then utilize our patch diffusion models to synthesize a large image.
@@ 4.1 Datasets
Figure 3: (Top) Patches (256 $\times$ 256) from our models, and the corresponding reference real patches used to generate them. The SSL-guided LDM replicates the semantics of the reference patch. (Bottom) Large images (1024 $\times$ 1024) from our models, and the corresponding reference real images used to generate them. We preserve the global arrangement of the semantics defined in the reference image.
@@ 4.2 Implementation details
We train our models on $6$ NVIDIA RTX 8000 GPUs, with a batch size of $100$ per GPU, utilizing code and pre-trained checkpoints from LDM [38]. We set the learning rate at $10^{-4}$ with a warmup of 10,000 steps. We apply DDIM with $50$ steps and a guidance scale of $1.75$ for both patch sampling and large-image generation. When generating large images we apply the patch diffusion model with a stride $>1$ depending on the desired target quality.
@@ 4.3 Image quality results
In Fig. 3, we present synthetic patches and large images from our TCGA-BRCA and NAIP models, along with the corresponding references from which the self-supervised embeddings were extracted. We evaluate our method’s per-patch and large-image generation quality by computing FID scores [14] using the Clean-FID implementation [32]. We generate 10,000 patches ($256\times 256$) and 3,000 large images ($1024\times 1024$) from diffusion models trained on the TCGA-BRCA, TCGA-CRC, and NAIP datasets. Since our generative model requires a self-supervised conditioning vector $y$ (or multiple vectors $z$) for each synthetic image (or large image), we randomly sample embeddings from reference images in the training set to generate images for evaluation.
For patches, we measure FID against the real images (“Vanilla FID"). For large images, we follow the evaluation strategy of MultiDiffusion [3] and use FID to compare the distribution of 256 $\times$ 256 crops from synthesized large images to that of real image patches of the same size (“Crop FID"). We also measure FID directly between the large images and ground truth data using CLIP [35] (“CLIP FID").
To evaluate the similarity between the reference and generated large images, we resize from $1024\times 1024$ to $256\times 256$ and use an SSL model to extract embeddings. We compute the cosine similarity between the paired embeddings (“Embedding Similarity").
Patch-level quality: Our models achieve low patch FID scores across all datasets (Tab. 1). For TCGA-BRCA patches, our “Vanilla FID" of 6.98 is comparable to the current state-of-the-art [49] ($7.64$ at 10×10$\times$10 ×). We attain similar, low FID scores for the smaller CRC and NAIP datasets.
@@
Table 1: FID scores for our generated patch and large images. Our patch-level BRCA model is on par with SoTA [49] (7.64 at 10 $\times$). “CLIP FID" and “Embedding Similarity" demonstrate our large images’ realism and contextual accuracy.
Synthetic large image quality: We present the “Crop FID", “CLIP FID" and “Embedding Similarity" results for large images in Table 1. The “Crop FID" of the large images is comparable to the “Vanilla FID" on the BRCA and CRC datasets, showing that patches from synthesized large images have similar semantic content to the ground truth patches. We attribute the worse “Crop FID" for the NAIP model to the limited number of samples available for both SSL and diffusion model training. The “Crop FID" is consistently higher than the patch-level FID, which is expected as the large-image generation framework only approximates the distribution of the large images and does not have access to the true conditioning at every location.
Our low “CLIP FID" scores indicate that the generated large images are similar to real images when resized to $224\times 224$ pixels. This indicates that our SSL-guided large-image generation successfully retains the larger-scale semantic arrangements of real data. For NAIP, our model is not as good at synthesizing high-frequency details, which explains the large discrepancy between “Crop FID" and “CLIP FID". Additionally, when comparing the “CLIP FID" of large images synthesized by the 20×20$\times$20 × BRCA model and resized to 5×5$\times$5 ×, to images from a model trained directly on 5×5$\times$5 × data, we see minimal difference (7.43 vs. 6.64).
We evaluate the contextual similarity between synthetic and reference large images for the BRCA and CRC data. We compute cosine similarity between large images using CTransPath embeddings. Our BRCA and CRC models demonstrate high “Embedding Similarity" scores of $0.924$ and $0.938$, respectively, reflecting our framework’s effectiveness in preserving the integrity of context and key features on the large-image scale.
@@ 5 Image Augmentation Experiments
As shown in Fig. 3, apart from visual fidelity, the synthetic images preserve the intricate characteristics of the reference images, both on the patch and large-image scale. These characteristics include the nuanced textural elements, histological staining, and cell structure. This correspondence between real and synthetic images demonstrates the detailed and varied information captured by the self-supervised embeddings used as conditioning. In conjunction with the high “Embedding Similarity", it justifies using variations of images generated from SSL embeddings for patch and large-image level data augmentation.
@@
Table 2: The inclusion of synthetic data consistently enhances AUC across various MIL architectures and BRCA tasks. The dataset contains 1000 real images, so “10% + synthetic" indicates training with 100 real and 100 synthetic WSIs, with the remainder used for testing.
@@
The results, presented in Table 3(a), reveal a notable improvement in classifier performance, from $78\%$ to $83\%$. This improvement again confirms the high quality of our synthetic data while also highlighting the versatility of our apporach. Despite being trained on $256\times 256$ patches from TCGA-BRCA, the model generalizes to produce realistic large images from a completely different dataset. We attribute this generalization capability to the expressiveness of the SSL features and the potency of the diffusion model to accurately portray them in images. While our model does not reach the current SoTA accuracy of $87\%$, achieved by an ensemble approach [7], the simplicity and ability to integrate with other models make for a noteworthy contribution.
@@
Table 3: Our data augmentations provide notable improvements for the BACH (a) and CRC-VAL-HE (b) datasets. Notably, the diffusion training data does not overlap with the data of the augmented datasets.
@@
We train an ImageNet pre-trained ResNet-50 [12] network on three splits: real images only, synthetic images only, and a combination of both. Evaluation on the CRC-VAL-HE-7K test set demonstrates a significant performance increase when synthetic data is introduced, with classifier accuracy rising from 93.8% to 96.27% as presented in Table 3(b). We match the current SoTA [20], which used an ensemble of deep models, with a model agnostic approach. As in the previous experiment, the diffusion model, now trained on the significantly smaller TCGA-CRC data, can also effectively synthesize images from a completely different dataset by only controlling the self-supervised conditioning.
@@ 7 Combining self-supervised embeddings with diffusion
We further evaluate our patch-generating models by posing the question does the diffusion model learn more about the data than the self-supervised learning model? We hypothesize that by combining the pre-trained self-supervised embedding with the denoising task we can improve the learned representations of the data, leading to better performance in downstream tasks, which are performed with features from self-supervised learning. To validate this hypothesis, we utilize the trained diffusion model as a feature extractor and apply a Multiple-instance learning (MIL) approach for the slide-level classification task of subtyping Breast Cancer. For each patch in a whole-slide image (WSI) we first extract the self-supervised embeddings, which are then used as conditioning to obtain features from the denoiser’s U-Net bottleneck layer at a fixed timestep $t=50$.
@@ Acknowledgements
This research was partially supported by NCI awards 5U24CA215109, 1R21CA258493-01A1, UH3CA225021, NSF grants IIS-2123920, IIS-2212046 and Stony Brook Profund 2022 seed funding. We thank Rajarsi Gupta for his valuable feedback.
@@ 9 Annotation costs
We asked a pathologist to annotate patches from TCGA-BRCA to estimate the cost of detailed per-patch annotation for the entirety of the dataset. We presented the 20×20$\times$20 × magnification patches of Fig. 5 and requested them to “write a brief description for each of the following patches". An expert pathologist required approximately 5-10 seconds to identify features and describe the patches. Therefore, for the entire 15M patches of TCGA-BRCA, it would take $\approx 40000$ hours to provide full per-patch annotations. This training dataset is small compared to the volume of data used in large studies, e.g. 10k whole slide images or approximately 10×10$\times$10 × the number of TCGA-BRCA data. Employing expert pathologists to annotate these vast amounts of data is prohibitively expensive and, therefore, practically infeasible at the scale at which we want to apply these models.